9. Given the following data, determine the nu- merical value of each of the expressions below. Person (i) X 9 8l 5 25 4 6 3le 3 4 d. EY c.ΣΧ, i 3 i-1
Answer #1
Person Xi Yi Xi +Yi XiYi
1 6 9 15 54
2 4 5 9 20
3 6 3 9 18
4 0 7 7 0
5 7 8 15 56
Total 23 32 55 148

a) ΣΧί = 23

b) y=945+3=17

c)Xi =640+7=13

d) 5

e) (X + Y) = 〉 (Xi Yi)-55

f)ΣΧί * Yi = 148